OpenStudy (anonymous):

What fraction is the same as 62/8? Not Here 7 5/8 7 3/8 7 1/8 8 1/4

OpenStudy (anonymous):

i got 7 3/4 so not here right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

62/8=7 +6/4=7 +3/4 so you are correct

OpenStudy (anonymous):

still stuck on that last average problem Andras

OpenStudy (anonymous):

so the average is the sum of the 5 numbers /5 so let x be the 5th number

OpenStudy (anonymous):

I will continue in the other topic :-)